radomes.openskiesproject.org
amrivera77
amrivera77 amrivera77
  • 17-01-2017
  • Mathematics
Answered

Use decimals to express a probability in tenths.


A shelf has 20 bags of potatoes. Two of the bags have potatoes starting to go bad.


What is the expected probability of the complement of the event "bad bags of potatoes"?

Verified Answer :

Annabri594 Annabri594
  • 17-01-2017

There are 2 bad bags out of 20, so the probability of bad bags is 2/10=1/10=0.1

Answer Link
lt974525 lt974525
  • 26-09-2018

A shelf has 20 bags of potatoes. Two of the bags have potatoes starting to go bad.

What is the expected probability of the complement of the event "bad bags of potatoes"?

A.0.1

B.0.2

C.0.8

D.0.9

Answer Link

Related Questions

How to calculate kinetic energy of a projectile at the highest point of a projectile
What was the long term effect of the Bleeding Kansas problem
Which Republican opposed the Dred Scott decision ACharles Sumner BRoger b Taney CJames Buchanan DMillard Fillmore
HELP A tree is an example of an Achemotroph B Scavenger C autotroph D heterotroph
You can use mole ratios and molar mass to determine the masses of substances involved in a chemical reaction. Using the periodic table, calculate the molar mass
Phosphates in a sentence
Find x intercepts for x squared 4x50
can someone help please with math
how too get to 104 only using 2358 once
Read the sentence below Simplicity is to garden design as symmetry is to building design What type of figurative language does this sentence use Analogy Allusio